OpenStudy (anonymous):

The slope of a line is 20 and the point (3, 5) is on the line. Express the line in slope-intercept form.

OpenStudy (anonymous):

y-5 = 20(x-3) y= 20x -60 +5 y=20x-55

OpenStudy (anonymous):

if i keep asking you questions will u answer them